# Softening holographic nuclear matter

### v1

@@ 1 Introduction
Matter at large baryon densities, for instance in the interior of neutron stars, is notoriously difficult to describe theoretically. Effective theories of weakly coupled nuclear matter Hebeler:2009iv ; Drischler:2016djf ; Tews:2018kmu are valid at comparably low densities and can at best be extrapolated to ultra-dense matter. Conversely, perturbative first-principle calculations from Quantum Chromodynamics (QCD) Freedman:1976ub ; Pisarski:1999tv ; Kurkela:2009gj are valid at asymptotically large densities where the QCD coupling constant is small Gross:1973id ; Politzer:1973fx , and one relies on extrapolations down in density. Together with results from numerous phenomenological models and extrapolations agnostic to the microscopic physics Annala:2017llu ; Altiparmak:2022bke , these approaches yield valuable insight into the condensed matter physics of QCD, but are far from providing a complete picture.
@@
Many-baryon systems require, in principle, many-instanton solutions in curved space. This is very difficult to handle, both analytically and numerically, and thus various degrees of simplifications have been employed, for instance to assume the instantons to be pointlike objects Bergman:2007wp . A crystalline structure in the spatial directions is expected for nuclear matter at large $N_{c}$ Klebanov:1985qi ; Rho:2009ym ; Kaplunovsky:2012gb ; Kaplunovsky:2013iza , and, additionally, the multi-instanton configurations can be expected to spread out in the holographic direction, possibly creating multiple “layers” as the density is increased Kaplunovsky:2012gb ; Preis:2016fsp ; Elliot-Ripley:2016uwb . At high densities, the instantons start to overlap, hence assuming a homogeneous configuration in the spatial directions becomes a good approximation. Real-world nuclear matter is homogeneous, at least around saturation density, which is experimentally accessible in large nuclei. Therefore, the homogeneous approximation can also be viewed as a (somewhat phenomenological) model approach, which might be closer to actual nuclear matter than the crystalline solutions that are plagued by large-$N_{c}$ artifacts.
@@ 2.1 Action
We consider the confined geometry of the Witten-Sakai-Sugimoto model with antipodal separation of the flavor branes, such that their embedding in the background geometry is geodesic and fixed. Moreover, we use the Yang-Mills (YM) approximation for the SU(2) gauge theory on the flavor branes, and include a Chern-Simons (CS) contribution that accounts for the topological baryon number. We write the gauge field action as
@@
Here, prime denotes derivative with respect to $u$. Within our homogeneous ansatz for isospin-symmetric matter it is consistent with the equations of motion to restrict ourselves to two independent nonzero components of the gauge field $A_{\mu}=\hat{A}_{\mu}{\bf 1}+A_{\mu}^{a}\sigma_{a}$ (with $\mu=0,1,2,3,u$ and the Pauli matrices $\sigma_{a}$, $a=1,2,3$): the temporal abelian component, which, for notational convenience, we denote by
@@
$A_{i}^{a}(u)=-\frac{\lambda_{0}}{2}h(u)\delta_{i}^{a}\,.$ (9)
@@
In all configurations we consider, we employ the boundary conditions
@@
The non-abelian gauge fields vanish at the boundary because a nonzero boundary condition for $h$ would lead to a divergent thermodynamic potential.
@@ 2.3 Equations of motion and free energy
The segment $u>u_{d}$ is different because it contains the chemical potential as a boundary condition and will thus be treated separately. These conditions are only relevant for 3 and 4 jumps, because for 1 or 2 jumps there is no segment between $u_{1}$ (or $u_{\rm KK}$) and $u_{d}$.
@@
The integral is finite and does not require any renormalization. It is normalized such that the vacuum solution $h(u)=0$ corresponds to $\Omega=0$.
@@ 2.5 Behavior at the discontinuities
Figure 2: Sketch and terminology for the 4 different configurations at a discontinuity that correspond to stationary points of the free energy. The lines illustrate the non-abelian gauge field in the form of the function $h(z)$ with $z$ increasing in each panel from left to right.
@@ 3.2 $D_{L}$ phase
Figure 5: Properties of phases with one (black) and two (blue) discontinuities, all at $\lambda=7.09$: Location of the discontinuity $u_{1}$ on one half of the flavor branes (upper left panel), baryon density $\bar{n}_{B}$ (upper right panel), and free energy density $\Omega$ (lower panel), all as functions of the chemical potential $\bar{\mu}_{B}$. Dots indicate the points of the onset points from the vacuum (and the point where $D_{R}$ connects to $D^{0}$). The $D_{A}$ configuration only exists at very large $\bar{\mu}_{B}$ (see upper left panel) and is outside the shown range of the two other panels. The vertical dashed lines in the upper panels indicate the first-order transitions between the vacuum and the $D^{0}$ phase (right panel) and between the $D^{0}$ and $D_{L}$ phases (both panels).
The existence of the $D_{L}$ phase was already pointed out (not in this terminology and without comparison to the $D^{0}$ phase) in ref. Rozali:2007rx . Since the holographic direction corresponds to an energy scale, the sharp boundary of the charge distribution was interpreted as a Fermi energy, possibly even representing a quark Fermi surface. We have not found any observable that would support this interpretation. In particular, it was later pointed out that quark degrees of freedom can be added in the form of string sources, giving rise to holographic quarkyonic matter (in combination with baryons in the pointlike approximation) Kovensky:2020xif , and thus it seems natural to interpret all baryon number in the approach of this paper to originate from baryons, not quarks [^3] We have checked within the deconfined geometry that string sources can also be added to the $D_{R}$ configuration, thus constructing a generalization of the quarkyonic phase of ref. Kovensky:2020xif .. At large $N_{c}$, baryons may show fermionic as well as bosonic properties and it might be tempting to speculate whether the different configurations we find here correspond to different fermionic and bosonic incarnations of large-$N_{c}$ nuclear matter. However, we have not found any consistent picture that would suggest an interpretation along these lines – for instance interpreting certain configurations as Fermi surfaces and others as Bose-Einstein condensates – and leave this aspect to future work.
@@ 3.5 $D_{A}$ phase
Let us use this example to explain the practical calculation in some more detail, to be applied to the other cases analogously. The system of equations can be decoupled to simplify the calculation. We can work with a given $z_{1}$ and first solve the differential equation for $h_{L}(z)$ with initial conditions (66a). This requires to find the value of $h_{(1)}$ such that the two first conditions of eq. (67) are met. In these conditions, $h_{1-}$ can be eliminated by inserting $h_{L}(z_{1})$ directly into the condition of the derivative. Then, $h_{1-}$ is simply extracted from the solution found for $h_{L}(z)$. Once we know $h_{1-}$, we can employ the equation of motion for $h_{R}(z)$. Recall that $h_{1-}$ enters this differential equation via the function $Q$ (21); conversely, $h_{1+}$ does not enter $Q$ in the equation of motion for $h_{L}(z)$. We determine $h_{1+}$ such that $h_{R}(z)$ with initial conditions (66b) has the correct boundary condition given in the third relation of eq. (67). The results are then used to compute $\bar{n}_{B}$ via eq. (65) and we obtain $\bar{\mu}_{B}$ and $\Omega$ in the usual way.
@@ 4.1 $D_{L}^{0}$ phase
In contrast to the phases with one or two jumps, the $D_{L}^{0}$ configuration has distinct classes of solutions. The reason is that $h_{L}(z)$ may or may not have a node, i.e., with $h_{c}<0$ it turns out that there are solutions for both $h_{1-}<0$ and $h_{1-}>0$. The profile of the former, which is the physically more interesting solution for reasons to be explained momentarily, is shown in the upper left panel of figure 7. (The same structure of solutions exists for many other configurations; the figure does show both solutions – with and without node – for the $D_{S}^{0}$ configuration.) We see that although we have two discontinuities in the gauge field, the charge distribution only has a single maximum, which we interpret as a single instanton layer. Just like the $D^{0}$ phase, and as can be anticipated from the barely visible shoulder-like structure in $\rho(z)$, the $D^{0}_{L}$ phase also turns dynamically into a 3-layer phase at large densities.
@@ 4.2 $D_{RL}$ phase
Now, for given $z_{1}$, we can solve the system of equations for the function $h(z)$ and the three variables $h_{1+}$, $h_{2-}$, $z_{2}$. Since this phase is trivial in the ultraviolet (like all phases whose sequence of jumps ends with an $L$-type jump), the chemical potential has the simple form (56) with $u_{1}$ and $z_{1}$ replaced by $u_{2}$ and $z_{2}$.
@@ 4.4 Free energy comparison and incompressibility
Figure 10: Left panel: Dimensionless incompressibility for the energetically most preferred 1-layer (black, $D^{0}$ and $D_{L}^{0}$) and 2-layer (blue, $D_{L}$ and $D_{RL}$) phases, as a function of baryon density, for $\lambda=7.09$. For the phases which have a first-order baryon onset, $D^{0}$ and $D_{L}$, the onset point is marked by a dot. The dashed lines are the results for the pointlike 1-layer (black), 2-layer (blue), and $\infty$-layer (magenta) phases. The energetically favored phase is $P_{\infty}$, closely followed by $D_{RL}$, see figure 9. Right panel: Incompressibility of the $D^{0}$ and $D_{RL}$ phases, computed with $M_{rmKK}=1000\,{\rm MeV}$, as a function of baryon density in units of the physical saturation density $n_{0}$. The background shows, for comparison, the incompressibility (of beta equilibrated, charge neutral matter) computed from various non-holographic models of nuclear matter (BL, DD2, LS220, SFHo), with asterisks indicating the central point of the heaviest stable star for each model; taken from ref. Perego:2021mkd , where also references for the individual models can be found.
@@ 5 Discussion
Finally, let us also discuss the significance of the pointlike solutions. In particular, one might argue that the continuum “2-layer” limit $P_{\infty}$ is, in some sense, the main result of the paper because this phase is energetically favored over all our multi-jump solutions. This observation is remarkable but has to be treated with care for several reasons. Firstly, we cannot claim to have done an exhaustive or even systematic study of multi-jump phases with more than 4 discontinuities of the non-abelian gauge field. More importantly, the pointlike approximation contains a singular ansatz without a corresponding solution of the equations of motion. Therefore, the $P_{\infty}$ phase might provide important clues on how to construct an actual solution on a more fundamental level, but its comparison to our multi-jump phases that do solve the equations of motion is, albeit instructive, not conclusive. Despite the different ansatz, it is striking that the charge distribution of the $P_{\infty}$ phase looks very similar to that of some of our multi-jump configurations and one may ask whether we have missed a certain solution that reproduces the $P_{\infty}$ phase. However, as just argued, there is no guarantee that the pointlike ansatz reproduces any actual solutions to the equations of motion. It is thus not surprising that even its continuum limit, where the charge distribution becomes piecewise smooth, yields a configuration that cannot be reproduced within our ansatz on the level of the gauge fields. Nevertheless, the novel $P_{\infty}$ phase is interesting for future applications as well, not only because of its low free energy but also due to its mathematical simplicity, which reduces the numerical effort to solving a simple algebraic equation. On the other hand, since it is not built upon an actual solution of the equations of motion it seems not straightforward to compute, e.g., transport properties in this phase.
@@ A.2 Continuum limit
To calculate the continuum limit, we first insert eq. (90b) into eq. (90a)

### v2

@@ 1 Introduction
Matter at large baryon densities, for instance in the interior of neutron stars, is notoriously difficult to describe theoretically. Effective theories of weakly coupled nuclear matter Hebeler:2009iv ; Drischler:2016djf ; Tews:2018kmu are valid at comparably low densities and can at best be extrapolated to ultra-dense matter. Conversely, perturbative first-principle calculations from Quantum Chromodynamics (QCD) Freedman:1976ub ; Pisarski:1999tv ; Kurkela:2009gj are valid at asymptotically large densities where the QCD coupling constant is small Gross:1973id ; Politzer:1973fx , and one relies on extrapolations down in density. Together with results from numerous phenomenological models and interpolations agnostic to the microscopic physics Annala:2017llu ; Altiparmak:2022bke , these approaches yield valuable insight into the condensed matter physics of QCD, but are far from providing a complete picture.
@@
Many-baryon systems require, in principle, many-instanton solutions in curved space. This is very difficult to handle, both analytically and numerically, and thus various degrees of simplifications have been employed, for instance to assume the instantons to be pointlike objects Bergman:2007wp . A crystalline structure in the spatial directions is expected for nuclear matter at large $N_{c}$ Klebanov:1985qi ; Rho:2009ym ; Kaplunovsky:2012gb ; Kaplunovsky:2013iza , and, additionally, the multi-instanton configurations can be expected to spread out in the holographic direction, possibly creating multiple “layers” as the density is increased Kaplunovsky:2012gb ; Preis:2016fsp ; Elliot-Ripley:2016uwb . At high densities, the instantons start to overlap, hence assuming a homogeneous configuration in the spatial directions becomes a good approximation. Real-world nuclear matter is homogeneous (a liquid, not a solid), at least around saturation density, which is experimentally accessible in large nuclei Reinhard:2005nj ; Horowitz:2020evx . Therefore, the homogeneous approximation can also be viewed as a (somewhat phenomenological) model approach, which might be closer to actual nuclear matter than the crystalline solutions that are plagued by large-$N_{c}$ artifacts.
@@ 2.1 Action
We consider the confined geometry of the Witten-Sakai-Sugimoto model with antipodal separation of the flavor branes, and we assume that their embedding in the background geometry is geodesic and fixed [^1] This is a stable configuration, at least in the absence of gauge fields Sakai:2004cn . We do not see an obvious reason why the baryonic solutions studied in this paper should change this. To prove stability of the probe branes rigorously, however, fluctuations on the background of the numerical gauge field solutions developed here must be considered. A full stability investigation (including other potential instabilities of our solutions, e.g., with respect to spatial inhomogeneities) is beyond the scope of this paper.. Moreover, we use the Yang-Mills (YM) approximation for the SU(2) gauge theory on the flavor branes, and include a Chern-Simons (CS) contribution that accounts for the topological baryon number. We write the gauge field action as
@@
Here, prime denotes derivative with respect to $u$. As can be seen from the general derivation Kovensky:2023mye , within our homogeneous ansatz for isospin-symmetric matter it is consistent with the equations of motion to restrict ourselves to two independent nonzero components of the gauge field $A_{\mu}=\hat{A}_{\mu}{\bf 1}+A_{\mu}^{a}\sigma_{a}$ (with $\mu=0,1,2,3,u$ and the Pauli matrices $\sigma_{a}$, $a=1,2,3$): the temporal abelian component, which, for notational convenience, we denote by
@@
$A_{i}^{a}(u)=-\frac{\lambda_{0}}{2}h(u)\delta_{i}^{a}\,$, (9)
@@
In all configurations we consider, we employ the boundary conditions Rozali:2007rx ; Li:2015uea
@@ 2.3 Equations of motion and free energy
The segment $u>u_{d}$ is different because it contains the chemical potential as a boundary condition and will thus be treated separately. For the configurations we consider later, these conditions are only relevant for 3 and 4 jumps, because for 1 or 2 jumps there is no segment between $u_{1}$ (or $u_{\rm KK}$) and $u_{d}$.
@@
The integral is finite and does not require any renormalization (since we have started from the YM action on the flavor branes, purely geometric contributions that are the same for all phases we consider are not part of $\Omega$). The vacuum solution $h(u)=0$ corresponds to $\Omega=0$.
@@ 2.5 Behavior at the discontinuities
Figure 2: Sketch and terminology for the 4 different configurations at a discontinuity that correspond to stationary points of the free energy. The lines illustrate the non-abelian gauge field in the form of the function $h(z)$ with $z>0$ increasing in each panel from left to right.
@@ 3.2 $D_{L}$ phase
Figure 5: Properties of phases with one (black) and two (blue) discontinuities, all at $\lambda=7.09$: Location of the discontinuity $u_{1}$ on one half of the flavor branes (upper left panel), baryon density $\bar{n}_{B}$ (upper right panel), and free energy density $\Omega$ (lower panel), all as functions of the chemical potential $\bar{\mu}_{B}$. Dots indicate the points of the onset from the vacuum (and the point where $D_{R}$ connects to $D^{0}$). The $D_{A}$ configuration only exists at very large $\bar{\mu}_{B}$ (see upper left panel) and is outside the shown range of the two other panels. The vertical dashed lines in the upper panels indicate the first-order transitions between the vacuum and the $D^{0}$ phase (right panel) and between the $D^{0}$ and $D_{L}$ phases (both panels).
The existence of the $D_{L}$ phase was already pointed out (not in this terminology and without comparison to the $D^{0}$ phase) in ref. Rozali:2007rx . Since the holographic direction corresponds to an energy scale, the sharp boundary of the charge distribution was interpreted as a Fermi energy, possibly even representing a quark Fermi surface. We have not found any observable that would support this interpretation. In particular, it was later pointed out that quark degrees of freedom can be added in the form of string sources, giving rise to holographic quarkyonic matter (in combination with baryons in the pointlike approximation) Kovensky:2020xif , and thus it seems natural to interpret all baryon number in the approach of this paper to originate from baryons, not quarks [^4] We have checked within the deconfined geometry that string sources can also be added to the $D_{R}$ configuration, thus constructing a generalization of the quarkyonic phase of ref. Kovensky:2020xif .. At large $N_{c}$, baryons may show fermionic as well as bosonic properties and it might be tempting to speculate whether the different configurations we find here correspond to different fermionic and bosonic incarnations of large-$N_{c}$ nuclear matter. However, we have not found any consistent picture that would suggest an interpretation along these lines – for instance interpreting certain configurations as Fermi surfaces and others as Bose-Einstein condensates – and leave this aspect to future work. Possibly the calculation of transport properties can shed some light on the physical interpretation.
@@ 3.5 $D_{A}$ phase
Let us use this example to explain the practical calculation in some more detail, to be applied to the other cases analogously. The system of equations can be decoupled to simplify the calculation. We can work with a given $z_{1}$ and first solve the differential equation for $h_{L}(z)$ with initial conditions (66a). This requires to find the value of $h_{(1)}$ such that the two first conditions of eq. (67) are met. In these conditions, $h_{1-}$ can be eliminated by inserting $h_{L}(z_{1})$ directly into the condition of the derivative. Therefore, the first step is reduced to determining the single variable $h_{(1)}$, which makes it easy to ensure that all possible solutions are found. Then, $h_{1-}$ is simply extracted from the solution found for $h_{L}(z)$. Once we know $h_{1-}$, we can employ the equation of motion for $h_{R}(z)$. Recall that $h_{1-}$ enters this differential equation via the function $Q$ (21); conversely, $h_{1+}$ does not enter $Q$ in the equation of motion for $h_{L}(z)$. We determine $h_{1+}$ such that $h_{R}(z)$ with initial conditions (66b) has the correct boundary condition given in the third relation of eq. (67). Again, this is a shooting problem with a single unknown variable. The results are then used to compute $\bar{n}_{B}$ via eq. (65) and we obtain $\bar{\mu}_{B}$ and $\Omega$ in the usual way.
@@ 4.1 $D_{L}^{0}$ phase
In contrast to the phases with one or two jumps, the $D_{L}^{0}$ configuration has distinct classes of solutions. The reason is that $h_{L}(z)$ may or may not have a node, i.e., with $h_{c}<0$ it turns out that there are solutions for both $h_{1-}<0$ and $h_{1-}>0$. The profile of the former, which is the physically more interesting solution for reasons to be explained momentarily, is shown in the upper left panel of figure 7. (The same structure of solutions exists for many other configurations; the figure does show both solutions – with and without node – for the $D_{S}^{0}$ configuration.) We see that although we have three discontinuities in the gauge field, the charge distribution only has a single maximum, which we interpret as a single instanton layer. Just like the $D^{0}$ phase, and as can be anticipated from the barely visible shoulder-like structure in $\rho(z)$, the $D^{0}_{L}$ phase also turns dynamically into a 3-layer phase at large densities.
@@ 4.2 $D_{RL}$ phase
Now, for given $z_{1}$, we can solve the system of equations for the function $h(z)$ and the three variables $h_{1+}$, $h_{2-}$, $z_{2}$. As in all cases we consider, and analogous to the details explained in the context of the $D_{A}$ phase, see Sec. 3.5, the system of equations can be decoupled. In this case it is two variables at most we have to solve for simultaneously, simple enough to systematically check for all possible solutions. Since this phase is trivial in the ultraviolet (like all phases whose sequence of jumps ends with an $L$-type jump), the chemical potential has the simple form (56) with $u_{1}$ and $z_{1}$ replaced by $u_{2}$ and $z_{2}$.
@@ 4.4 Free energy comparison and incompressibility
Figure 10: Left panel: Dimensionless incompressibility for the energetically most preferred 1-layer (black, $D^{0}$ and $D_{L}^{0}$) and 2-layer (blue, $D_{L}$ and $D_{RL}$) phases, as a function of baryon density, for $\lambda=7.09$. For the phases which have a first-order baryon onset, $D^{0}$ and $D_{L}$, the onset point is marked by a dot. The dashed lines are the results for the pointlike 1-layer (black), 2-layer (blue), and $\infty$-layer (magenta) phases. The energetically favored phase is $P_{\infty}$, closely followed by $D_{RL}$, see figure 9. Right panel: Incompressibility of the $D^{0}$ and $D_{RL}$ phases, computed with $M_{\rm KK}=1000\,{\rm MeV}$, as a function of baryon density in units of the physical saturation density $n_{0}$. The background shows, for comparison, the incompressibility (of beta equilibrated, charge neutral matter) computed from various non-holographic models of nuclear matter (BL, DD2, LS220, SFHo), with asterisks indicating the central point of the heaviest stable star for each model; taken from ref. Perego:2021mkd , where also references for the individual models can be found.
@@ 5 Discussion
Finally, let us also discuss the significance of the pointlike solutions. In particular, one might argue that the continuum “2-layer” limit $P_{\infty}$ is, in some sense, the main result of the paper because this phase is energetically favored over all our multi-jump solutions. This observation is remarkable but has to be treated with care for several reasons. Firstly, we cannot claim to have done an exhaustive or even systematic study of multi-jump phases with more than 4 discontinuities of the non-abelian gauge field. More importantly, the pointlike approximation contains a singular ansatz without a corresponding solution of the equations of motion, with a pointlike source whose only direct coupling is to the abelian gauge field. Therefore, the $P_{\infty}$ phase might provide important clues on how to construct an actual solution on a more fundamental level, but its comparison to our multi-jump phases that do solve the equations of motion is, albeit instructive, not conclusive. Despite the different ansatz, it is striking that the charge distribution of the $P_{\infty}$ phase looks very similar to that of some of our multi-jump configurations and one may ask whether we have missed a certain solution that reproduces the $P_{\infty}$ phase. However, as just argued, there is no guarantee that the pointlike ansatz reproduces any actual solutions to the equations of motion. It is thus not surprising that even its continuum limit, where the charge distribution becomes piecewise smooth, yields a configuration that cannot be reproduced within our ansatz on the level of the gauge fields. Nevertheless, the novel $P_{\infty}$ phase is interesting for future applications as well, not only because of its low free energy but also due to its mathematical simplicity, which reduces the numerical effort to solving a simple algebraic equation. On the other hand, since it is not built upon an actual solution of the equations of motion it seems not straightforward to compute, e.g., transport properties in this phase.
@@ A.2 Continuum limit
To calculate the continuum limit, we first insert eq. (90b) into eq. (90a),